A rectangular box with a square base is 11 inches wider than it is tall. If x represents the width of the base of the box, then the volume of the box as a function of x is

Answer by ankor@dixie-net.com(22740)

A rectangular box with a square base is 11 inches wider than it is tall. If x represents the width of the base of the box, then the volume of the box as a function of x is
:
x = the width
A square base therefore:
x = length
and "a square base is 11 inches wider than it is tall"
(x-11) = the height
:
f(x) = the volume
f(x) = x * x * (x-11)
f(x) = x^2(x-11)
f(x) = x^3 - 11x^2